Suppose Rocky Brands has earnings per share of ​$2.31 and EBITDA of ​$30.4 million. The firm also has 4.8 million shares outstanding and debt of ​$115 million​ (net of​ cash). You believe Deckers Outdoor Corporation is comparable to Rocky Brands in terms of its underlying​ business, but Deckers has no debt. If Deckers has a​ P/E of 13.1 and an enterprise value to EBITDA multiple of 7.1​, estimate the value of Rocky Brands stock using both multiples. Which estimate is likely to be more​ accurate?
Answer #1

a)

Value of 1 Share using P/E Ratio = EPS * P/E Ratio

=2.31*13.1

=30.261

Value of Company = Value of 1 share * outstanding shares

=30.261* 4.8 million

=145,252,800

b)

Enterprise value = EBITDA * EBITDA multiple

=30.4 million * 7.1

=215,840,000

Value of company = enterpise value - debt

=215,840,000 - 115,000,000

=100,840,000

EBITDA multipler is more likely to be accurate.